OpenStudy (anonymous):

Divide: 14x^7-6x^3/2x^3

OpenStudy (saifoo.khan):

\[\Large \frac{14x^7}{2x^3} - \frac{6x^3}{2x^3}\]Can you solve now?

OpenStudy (anonymous):

Is it 7x^4-3

OpenStudy (saifoo.khan):

Perfect.

OpenStudy (anonymous):

My answer was 7x^4-3 is that right?

OpenStudy (anonymous):

If this is the original equation, then yes: \[\Large{\frac{14x^{7}-6x^{3}}{2x^{3}}}\]

OpenStudy (anonymous):

Yeah it is, thanks!